OpenStudy (31356):

Quadratic Equations 6x=4-18x^2 MEDAL REWARDED!!

OpenStudy (31356):

@texaschic101 ?

OpenStudy (31356):

I know what the quadratic equation is :D

OpenStudy (31356):

I am just confused with this, that's all.

OpenStudy (31356):

@whpalmer4 ?

OpenStudy (anonymous):

the answer I got was 3

OpenStudy (chris911):

go here http://www.geteasysolution.com/6x=4-18x%5E2

OpenStudy (31356):

This is Quadratic Equation @Chris911, thanks anyways :D

OpenStudy (31356):

@MarziaJacksonCute, can you tell me how you got 3?

OpenStudy (31356):

I know the first thing we're supposed to do is to make it (ax^2+bx=c) & equal to 0

OpenStudy (anonymous):

Add '3' to each side of the equation. -3 + 3 + x2 = 0 + 3 Combine like terms: -3 + 3 = 0 0 + x2 = 0 + 3 x2 = 0 + 3 Combine like terms: 0 + 3 = 3 x2 = 3 Simplifying x2 = 3 ( that's how my teacher told me)

OpenStudy (31356):

Hmm.... I don't think this is Quadratic Equations? We are supposed to use \[x=-b +/- \sqrt{b^2 - 4ac} / 2a\]

OpenStudy (31356):

@MarziaJacksonCute ?

OpenStudy (anonymous):

I know but that's what my teacher told me

OpenStudy (31356):

Okay..............

OpenStudy (31356):

@mathslover ?

OpenStudy (31356):

Thx for the medal :D

mathslover (mathslover):

Well first try to bring all the terms on a same side : \(\mathsf{6x = 4 - 18x^2 }\\ \textbf{Transpose -18x^2 and 4 to the Left Hand Side} \\ \sf{6x - 4 + 18x^2 = 0 \\ 18x^2 +6x - 4 = 0 }\)

OpenStudy (31356):

Okay

mathslover (mathslover):

Now compare the above equation ( \(\sf{18x^2 + 6x - 4 =0 }\)) with the basic quadratic equation ( \(\sf{ax^2 + bx + c = 0}\) )and find the values of a , b and c. Can you do this?

OpenStudy (31356):

Thanks @mathslover I can do it from there, :D

OpenStudy (31356):

Can you help me with other similar problems?

OpenStudy (31356):

I know how to solve, just confused how to rearrange it that's all :D

mathslover (mathslover):

You're welcome! Good Luck and let me know what you get so that I can verify it. Also, I will try to help but as I told earlier that today is my exam so I may leave any time. Just tag me in your question, I will be there!

OpenStudy (31356):

Hmmm....... x^2=8x?

OpenStudy (31356):

Okay @mathslover, I wish you good luck on your exam! :D

OpenStudy (anonymous):

you can think of \[x^2 = 8x\] as \[x^2 - 8x + 0 = 0\] so what two numbers multiplied together are 0 but add to -8. Only two numbers are -8 and 0 so you write \[(x-8)(x+0) = 0\] hence x = 8 and 0

OpenStudy (31356):

And how about 4x^2=9?

mathslover (mathslover):

There are two methods to do this : i) Transpose 8x to Left Hand side : \(\sf{x^2 - 8x = 0}\) And now compare this with \(ax^2 + bx + c = 0\) as you can see, c = 0 (there is no constant term) and b is "-8" and a is "1" Just plug-in these values in the quadratic equation and solve : \(\color{blue}{\sf{\cfrac{-b \pm \sqrt{b^2-4ac}}{2a}}}\) ii) Divide both sides by "x" \(\color{orange}{\sf{\cfrac{x^2}{x}} }= \color{red}{\sf{\cfrac{-8x}{x}}}\) Cancel the possible terms and find x.

OpenStudy (31356):

Thx both of you! :D I wish I can give 2 medals though

mathslover (mathslover):

* it is 8x not "-8x" (in the second way)

OpenStudy (31356):

Okay

mathslover (mathslover):

Take my medal back and give it to Tyler @31356 - I have enough medals for now. Tyler explained well and I will be more happy if Tyler is awarded with a medal.

OpenStudy (31356):

Okay, lol :D

OpenStudy (31356):

There we go :D

OpenStudy (31356):

And how about 4x^2=9?

mathslover (mathslover):

I should go now. Bye! And Good Luck. Also, good work @Tyler1992 - Keep it up.

OpenStudy (31356):

And how about 4x^2=9?

mathslover (mathslover):

Take square root both sides. And solve for x. Or transpose 9 to LHS 4x^2 - 9 = 0 Solve it by finding the values of a, b and c and then put it in quadratic equation. Tyler will explain you how to solve it! I have to go. Bye for now.

OpenStudy (31356):

Okay thanks :D

OpenStudy (anonymous):

\[4x^2 = 9\] can be written as \[4x^2 + 0x - 9 = 0\] and then best to use quadratic formula i think.

OpenStudy (anonymous):

or i can show you another way to factor that if you want

OpenStudy (31356):

Oh it's fine :D Thanks though

OpenStudy (31356):

Thanks for your help :D I figured out the others :D

OpenStudy (anonymous):

No problem! :)

OpenStudy (31356):

Wait........

OpenStudy (31356):

How about 3x^2-10=12-31x?

OpenStudy (31356):

Now that one is a bit awkward?

OpenStudy (anonymous):

Its looks tricky but its actually not so bad :) so first you want to move everything on the right side to the left\[3x^2 - 10 -12 - 31x = 0\] now just combine like terms\[3x^2 - 31x - 22 = 0\] then use quadratic

OpenStudy (31356):

Okay, thanks! :D

OpenStudy (anonymous):

and i made a mistake

OpenStudy (31356):

Where?

OpenStudy (anonymous):

+31x

OpenStudy (31356):

Oh okay, thanks for the correction :D

OpenStudy (31356):

No more problems left :D

OpenStudy (31356):

Thanks! :D

OpenStudy (anonymous):

You're welcome!

OpenStudy (31356):

I can't stop saying "Thanks" lol :D

OpenStudy (31356):

Got to go, bye :D

OpenStudy (anonymous):

See ya :)